OpenStudy (anonymous):

SAT problem 18. The price of corn meal is d dollars for 12 ounces and each ounce makes c squares of cornbread. In terms of c and d, what is the dollar cost of the corn meal required to make 1 square of cornbread? (A) dc/12 (B) 12/dc (C) d/12c (D) c/12d (E) 12dc

OpenStudy (mertsj):

\[1sq \times \frac{1oz}{c squares}\times \frac{d dollars}{12 oz}= \frac{d}{12c} dollars\]